Sunday, April 23, 2006
Eric Update, Noon, 23 April 06
Eric much better today. All stats normal, ie blood pressure, heart rate, oxygen etc. Surgeon Mr M. visited and was pleased with Eric's progress. Eric eating and taking in lots of fluid. Still very weak and tired but seems to have turned the corner. More later. Lindsay. PS Family visitors only today and tomorrow. Please ring Lindsay (0207 640 2306) if you wish to visit.